OpenStudy (anonymous):

For what value(s) of the variable is the expression undefined? d-2/3d^2-10d+8

OpenStudy (callisto):

When it is undefined, the denominator =0 So, put 3d^2-10d+8 =0 (3d-4)(d-2)=0 (3d-4)=0 or (d-2)=0 d = 4/3 or d=2 So, when d =4/3 or d =2 , the expression is undefined~

OpenStudy (callisto):

Note that anything divided by 0 is undefined :|

OpenStudy (campbell_st):

factorise the denominator \[\frac{d -2}{(3d - 4)(d -2)}\] this simplifies to \[\frac{1}{3d - 4}\] there are a couple of things in the problem there restrictions on the domain are all real x except x cannot equal 2 or 4/3 x = 4/3 will be a vertical asymptote and x = 2 is a point of discontinuity